A pharmacist wishes to mix a solution that is 55​% Minoxidil. She has on hand 6060 ml of a 44​% solution and wishes to add some 77​% solution to obtain the desired 55​% solution. How much 77​% solution should she​ add?
Answer by jorel1380(3719)

Let n be the amount of 77% solution to be added. Then:
0.77n + 0.44(6060)=0.55(6060+n)
0.77n+2666.4=3333+0.55n
0.22n=666.6
n=3030 ml of 77% solution to be added
☺☺☺☺